A train leaves a city heading west and travels at 50 miles per hour. Three hours later, a second train leaves from the same place and travels in the same direction at 65 miles per hour. How long will it take for the second train to overtake the first train?
Answer by mananth(16949)

A train leaves a city heading west and travels at 50 miles per hour. Three hours later, a second train leaves from the same place and travels in the same direction at 65 miles per hour. How long will it take for the second train to overtake the first train?
..
d=rt
d= 50*3= 150 miles.
when the train B starts the train A has traveled 150 miles.
rate train A = 50
rate train B = 65
..
same direction
65-50 = 15 mph is the effective speed
Train B catches up 15 mph
distance =120
so 120/15 = 8 hours
...
m.ananth@hotmail.ca